OpenStudy (anonymous):

(-x+7)(2x) - 3(2x) = -1

OpenStudy (moongazer):

(2x)*(-x)+(2x)*7 - 3*2x =-1 solve it :)

OpenStudy (anonymous):

-2x+14x-6x=-1?

OpenStudy (moongazer):

-2x^2+14x-6x=-1

OpenStudy (anonymous):

how did you come up with an exponent?

OpenStudy (moongazer):

cause you multiplied (-x)*((2x)

OpenStudy (anonymous):

oh okay i see now.

OpenStudy (moongazer):

when you multiply same variables you add their exponents

OpenStudy (anonymous):

oh okay. Do I add like terms?

OpenStudy (moongazer):

yup

OpenStudy (anonymous):

so would i end up with: 6^2=-1?

OpenStudy (moongazer):

where did you get that ?

OpenStudy (anonymous):

i add the -2x^2 and 14x then subtracted 6.

OpenStudy (moongazer):

You can't add -2x^2 and 14x they have different variables

OpenStudy (anonymous):

okay then i would do 14x-6x=8x: 2x^2+8x=-1?

OpenStudy (anonymous):

im confused now...

OpenStudy (moongazer):

-2x^2+8x=-1 you are almost right you forgot the negative sign

OpenStudy (anonymous):

-2x^2+8x=-1

OpenStudy (anonymous):

typo. sorry

OpenStudy (anonymous):

i dont know what that means.. yet.

OpenStudy (moongazer):

transfer -1 to the right side of the equation :)

OpenStudy (anonymous):

how would i do that if there are both variables on the right and -1 isnt a variable?

OpenStudy (moongazer):

it is like adding both sides of the equation by +1 . so that you will cancel out -1 on the right side

OpenStudy (anonymous):

where would I add it since there both variables? could you show me an example?

OpenStudy (moongazer):

2x+1=2 2x+1-2=2-2 2x-1=0